# Dynamic Input Mapping Inversion for Algebraic Loop-Free Control in Hydraulic Actuators

### v2

@@ VI-A Standard controllers for study case
The control action for the standard cases are obtained by following the procedure described in Sec. III. The input gain is defined as
$g=\frac{lA_{p}B}{J_{tot}V_{p}}K_{Qp}+\frac{lA_{r}B}{J_{tot}V_{r}}K_{Qr}.$ (49)
The feedback linearization controller is derived as,
[…]
@@
where the angle dependency is removed, since it only affects the $r^{(3)}$ term and the term associated with the gain, which is a degree of freedom anyway.
The same applies for the sliding mode controller:
$u_{ST}=\frac{1}{g}$ $\left[\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}\right.$ (51)
$\,\,\left.-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s)d\tau\right],$
with
$\begin{bmatrix}\dot{e}\\ \ddot{e}\end{bmatrix}\approx\begin{bmatrix}v_{\beta}-\dot{r}\\ \displaystyle-\frac{\hat{M_{d}}}{J_{tot}}+\frac{l(A_{p}p_{p}-A_{r}p_{r})}{J_{tot}}-\ddot{r}\end{bmatrix}$ (52)
and $\hat{M_{d}}$ the new disturbance being estimated.
VI-B Mapping inversion based controllers for study case
@@
Following the same approach, the virtual control input in the proposed solutions are:
$v_{FL}=\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}$ (53)
$v_{ST}=$ $\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}$ (54)
$\,\,-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))d\tau,$
and the dynamical mapping inversion algorithm is,
@@
$\rho(t,\boldsymbol{x},u)=\frac{lA_{p}B}{J_{tot}V_{p}}Q_{p}(p_{p},p_{r},u)+\frac{lA_{r}B}{J_{tot}V_{r}}Q_{rv}(p_{r},u)$
@@
where $K(u)$ is defined according to Eq. 31 as
$K(u)=10^{-7}\left(-23.67u^{5}+40.73u^{3}+1.97u\right)\;.$ (56)
VII Simulation Results
The designed control actions are tested on a high-fidelity simulator, which includes the general dynamics presented in Eqs. 1,45, and also encompass the control valve dynamics. The measured signals are subjected to noise, simulated at realistic levels. The velocities ($v_{c}$ and $v_{\beta}$) are obtained via numerical differentiation since they are not usually available as sensor signals in a real plant. The pitch angle reference and the disturbance $M_{ext}$ are obtained from a high-fidelity simulator for a wind turbine based on OpenFAST that is widely used in wind turbine research and development.
Figure 6: Tracking performance of the Feedback linearization based controller, with and without applying the dynamical mapping inversion.
@@ VII-A Tracking performance
The tracking performance of the feedback linearization controller is presented in Fig. 6, and in Fig. 7 in a zoomed version. Three controllers are presented, for comparing the standard Feedback Linearizatrion (FL) solution with the one where the dynamic mapping inversion is applied (MI-FL) under the same tuning gains, and by overtuning (OT) the standard solution to be able to match the performance of the MI-FL controller. While both controllers are showing good tracking performance, the MI-FL (for a given tuning) shows a smaller error over the simulation window. By increasing the gains, it is possible to match its performance, with the consequence of getting an increase of unwanted chattering in the input.
@@
The zoomed plot gives a visualization of the input behaviour $u$, and of the valve position $x_{v}$. The MI-FL controller shows a reduced chatter in the regions where the pitch is varying in amplituide, while it shows a more chattery behaviour when the pitch is constant or slowly varying. Furthermore, the signal quality in $u$ and $x_{v}$ are equivalent for the MI solution, while part of the noise is filtered out by the valve dynamics in the standard solution, since the controller doesn’t have knowledge of the valve dynamics, nor of valve limitations.
The results for the ST-SMC are stated in Fig. 8.
Figure 8: Tracking performance of the Supertwisting sliding mode controllers, with and without applying dynamical mapping inversion.
The tracking performances are more similar between the two solutions with respect to the FL controller. The OT version doesn’t improve the tracking performance although leading to a more chattery input. The disturbance rate that the controller needs to handle, is shown in Fig. 9, where it is evident how higher gains lead to a larger amplitude in the command signal.
Figure 9: Perturbation derivative of sliding mode dynamics under different STSMC.
The unexpected dependency of the disturbance rate on the tuning gains is justified by the use of an observer for estimating the disturbance. Higher gains lead to a higher error in the estimation, which is affecting the overall control performance. The performance comparison of the observer with the different controllers is stated in Fig. 10. The tuning doesn’t coincide with the bounds since higher gains are reflected with higher disturbance estimation errors and, thus, higher disturbance derivative bounds. The STSMC is not capable of ensuring convergence on the origin at any time instant, but boundedness is guaranteed, according to [24].
@@ VII-B Mapping inversion algorithm
The tracking performance of the mapping inversion algorithm is presented in Fig. 11. $v$ and $\rho$ are the signal defined in Eq. 55, while $v_{sim}$ is the equivalent of the $v$ signal which is obtained in simulation, hence based on the real states of the system and not on the noisy measurements. The algorithm is not able to perfectly reconstruct $v$, because of the presence of noise in the signals. However, it is shown how it is possible to determine a proper tuning such that it is possible to reconstruct the actual virtual input in the plant ($v_{sim}$) and filtering out the noise present in the measurements.
Figure 10: Wind load disturbance estimation and comparison with LPF hydraulic force.
Thus, there is a natural filtering action which can be approximate with a simple first order low-pass filter in order to determine an indication of the bandwidth of such filtering action, in comparison to the bandwidth of the valve which represents a physical limitation in what the control action can deliver on the real system. The comparison is presented in Fig. 12.
Figure 11: Dynamical mapping inversion algorithm capability of tracking the required virtual controller, with a filtering action that allows to reconstruct the actual signal value in simulation
The cutoff frequency of the MI algorithm is lower than the valve’s bandwidth, meaning the valve only needs to follow trajectories it can actually withstand. With respect to the standard solution, where the filtering was occurring in the valve directly, in this new solution is the algorithm accounting for the valve limitations and filtering the input signal. This result support the similar signal quality in $u$ and $x_{v}$ in the MI solution that was found in Fig. 7.
Figure 12: Bode plots of valve dynamics and mapping inversion filtering effect.
@@
The tuning presented shows some room with respect to the maximum bandwidth that can be achieved by the valve. However, a faster tuning is reflected in a noisier reconstruction of $v_{sim}$ from the noisy measurements, hence more chattery $u$ and $x_{v}$. The slew rate limitation of the valve may also play a role.
@@ VII-C Controllers comparison
Assessing the tracking performance based on the error is straightforward, while assessing the signal quality of the input is more challenging since it is not a constant average signal. The solution was to calculate a moving average $\bar{u}$ and defining a new signal $u_{f}=u-\bar{u}$ which has zero mean and still maintains the information of the high frequency oscillation. The moving average window is the same for all the tests. The same applies for determining $x_{v,f}$. An example of a $u_{f}$ signal is illustrated in Fig. 13.
@@
  • $\text{RMS}(u_{f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}u_{fi}^{2}}$
  • $\text{RMS}(x_{v,f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{v,fi}^{2}}$
  • $\text{MAX}(e)=\max\limits_{0\leq t\leq T}e(t)$
  • $\text{RMS}(e)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}e_{i}^{2}}$
  • $\text{QWAE}(e)=\sqrt{\frac{\sum_{i=1}^{N}e_{i}^{2}|e_{i}|}{\sum_{i=1}^{N}e_{i}^{2}}}$
@@
where $T$ is the simulation time and the resulting values are normalized for having a clearer comparison.
Figure 13: Moving average removal for deriving a metrics associated with signal chatter.
The QWAE (Quadratically Weighted Average Error) metric is used to penalize larger errors more heavily, while the ISE (Integral Squared Error) metric measures the persistence of errors.
The result for the system running with the different controllers is shown in Fig. 14.
The solution proposed in the article performs better than standard one with the same tuning and with higher gains. The only exceptions are the cases of the $RMS(e)$ and the $ISE(e)$, where the OT ST-SMC displays a slightly smaller error, at the price of having much worse input performance. For a fixed tuning, the error with the MI is lower than when applying the respective standard algorithm.
The input signal is significantly improved by the filtering action of the MI, while the improvement in the actual valve position becomes less significant thanks to the filtering action of the valve itself. Still, the solution that leverages the MI shows the cleanest input over a window where both fast varying and constant references are required. The amplitude of $\text{RMS}(u_{f})$ and $\text{RMS}(x_{v,f})$ are equivalent in the MI case.
Figure 14: Bar plots for evaluating performance of the different controllers according to different metrics, associated with control input and tracking error.
@@ VIII Conclusion
This article presented a solution to eliminate the formation of an algebraic loop when designing some conventional nonlinear controllers. This was achieved by design of a virtual input and dynamically deriving a control action, which emulated the desired behaviour, while control valve trajectory commands were avoided if they would exceed the specified admissible range for valve behaviours.
The effectiveness of the proposed solution was tested in high-fidelity simulation for the industrial application of wind turbines’ blade pitch control, showing a general improvement in tracking performance and an essential reduction of input chatter, especially in the regions where the pitch reference is highly varying.
The dynamic inversion solution enabled room for flexibility and allows for future progress with other nonlinear parameter estimation algorithms.
A further step towards industrial implementation will be to derive the discrete-time version of the algorithm.
@@ Acknowledgements
This research was funded by The Danish Energy Technology Development and Demonstration Programm (EUDP) through the project: “Decreased Cost of Energy (CoE) from wind turbines by reducing pitch system faults”, grant number 64022-1058. The authors much appreciate this support.
[Feedback interconnected systems derivations] Let $\delta_{\underline{\mathbf{e}}}$ be the error associated with the calculation of $\ddot{e}$. The closed-loop system with the FL controller in the form of $\Sigma_{FL}$ is obtained as:
$\dot{\underline{\mathbf{e}}}=$ $\boldsymbol{A_{FL}}\underline{\mathbf{e}}+\boldsymbol{B_{FL}}(v^{*}-\tilde{v})$ (57)
$=$ $\boldsymbol{A_{FL}}\underline{\mathbf{e}}+\boldsymbol{B_{FL}}\left(-K_{FL}(\underline{\mathbf{e}}-\delta_{\underline{\mathbf{e}}})-\frac{\dot{\tilde{F}}_{d}}{M}-\tilde{v}\right)$
$=$ $(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{B_{FL}}\left(K_{FL}\begin{bmatrix}0\\ -\frac{\tilde{F}_{d}}{M}\end{bmatrix}-\frac{\dot{\tilde{F}}_{d}}{M}\right)$
$-\boldsymbol{B_{FL}}\,\left[\rho(\tilde{u})-\rho(u^{*}-\tilde{u})\right]$
$=$ $(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{\xi_{F_{d}}}+\boldsymbol{h_{FL}}(t,\tilde{u}).$
The unperturbed system is $\boldsymbol{h_{FL}}(t,0)=\boldsymbol{0}$ and the remaining terms describe the closed-loop dynamics with the standard design. Except for the bounded term $\boldsymbol{\xi_{F_{d}}}$, the system is exponentially stable with a convergence rate determined by the eigenvalues of $\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl}$. The bounded term acts as a disturbance that causes the system to be UUGB with sufficiently high gains.
Let
$\tilde{s}=s-s^{*},$ (58)
where $s$ is the actual value of the sliding variable and $s^{*}$ is calculated by differentiating the error with a robust exact differentiator, like the one in [25]. The sliding variable dynamics in closed loop reads:
$\dot{s}=$ $-k_{1}\lfloor s^{*}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s^{*}(\tau))d\tau$ (59)
$-k_{1}\lfloor\tilde{s}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(\tilde{s}(\tau))d\tau-\frac{\dot{\tilde{F}}_{d}}{M}-\lambda_{2}\frac{\tilde{F}_{d}}{M}$
$-\rho(u^{*}-\tilde{u})$
$=$ $-k_{1}\lfloor s^{*}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s^{*}(\tau))d\tau-\Delta(t)-\rho(u^{*}-\tilde{u}),$
where all the terms in $\Delta$ are bounded: $\tilde{s}$ is bounded and with zero mean, since the reconstruction error with the robust exact differentiator is only related to noise, and the disturbance estimations are bounded (proof in Sec. III-C1). For $s\neq 0$ one can define [22]
$\zeta_{1}\triangleq|s^{*}|^{1/2}sgn(s^{*})$ (60)
[…]
$\boldsymbol{\zeta}\triangleq\begin{bmatrix}\zeta_{1}&\zeta_{2}\end{bmatrix}^{T},$
and rewrite the closed-loop system as:
$\boldsymbol{\dot{\zeta}}=\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\displaystyle\frac{1}{2}k_{1}&1\\ -k_{2}&0\end{bmatrix}\boldsymbol{\zeta}+\begin{bmatrix}0\\ -d\end{bmatrix}+\underbrace{\begin{bmatrix}-\displaystyle\frac{\rho(u-\tilde{u})}{|\zeta_{1}|}\\ 0\end{bmatrix}}_{\displaystyle\boldsymbol{h}(t,\tilde{u})},$ (61)
with $k_{1},k_{2}$ selected to fulfil the conditions in [26]. Since $h(t,0)=0$, the unperturbed system has a finite stable equilibrium point at the origin. This implies (see [22]) the existence of a diffeomorphism $T:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ such that for $\boldsymbol{\vartheta}=T(\boldsymbol{\zeta})$ and
$\boldsymbol{\dot{\vartheta}}=\frac{\partial T}{\partial\boldsymbol{\zeta}}\left\{\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\frac{k_{1}}{2}&1\\ -k_{2}&0\end{bmatrix}T^{-1}(\boldsymbol{\vartheta})-\begin{bmatrix}0\\ d\end{bmatrix}\right\}-\frac{\partial T}{\partial\boldsymbol{\zeta}}\begin{bmatrix}\displaystyle\frac{1}{|\zeta_{1}|}\\ 0\end{bmatrix}\rho(\tilde{u}-\hat{u}),$ (62)
$\exists V_{ST}=\boldsymbol{\vartheta}^{T}P_{ST}\,\boldsymbol{\vartheta}$ such that $V_{2}\leq-c||\boldsymbol{\vartheta}||\triangleq-W_{ST}(||\boldsymbol{\vartheta}||)$.

### v4

@@ VI-C1 Standard controller
The control action for the standard case is obtained by defining the input gain:
$g=\frac{r_{\beta}A_{p}B}{J_{tot}V_{p}}K_{Qp}+\frac{r_{\beta}A_{r}B}{J_{tot}V_{r}}K_{Qr}.$ (43)
Finally, the feedback linearization controller is derived as,
$\begin{aligned} u&=\frac{1}{g}\left[\frac{r_{\beta}A_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{r_{\beta}A_{r}^{2}B}{J_{tot}V_{r}}v_{c}+\frac{\dot{\hat{M}}_{d}/J_{tot}+r^{(3)}-K_{FL}\underline{\mathbf{e}}_{\beta}}{\sin({\psi_{\beta}(x_{c})})}\right]\\ &\approx\frac{1}{g}\left[\frac{r_{\beta}A_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{r_{\beta}A_{r}^{2}B}{J_{tot}V_{r}}v_{c}+\frac{\dot{\hat{M}}_{d}}{J_{tot}}+r^{(3)}-K_{FL}\underline{\mathbf{e}}_{\beta}\right]\end{aligned},$ (44)
@@
where $\hat{M_{d}}$ is available from the load cylinder, and the $\psi_{\beta}$ angle dependency is removed since it only affects the $r^{(3)}$ term and the term associated with the gain, which is a degree of freedom anyway. The vector $\underline{\mathbf{e}}_{\beta}$ contains the tracking error $e_{\beta}$ and its derivatives, similarly to (10).
Figure 9: Schematics of the experimental setup with pitch control cylinder right, generating $F_{h}$, and load cylinder left, generating $F_{h,L}$ in real time.
VI-C2 Mapping inversion based controller
@@
Following a similar approach, the virtual control input in the proposed solution is:
$v=\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{r_{\beta}A_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{r_{\beta}A_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}_{\beta},$ (45)
and the dynamical mapping inversion algorithm is
@@
$\rho(t,\boldsymbol{x},u)=\frac{r_{\beta}A_{p}B}{J_{tot}V_{p}}Q_{p}(p_{p},p_{r},u)+\frac{r_{\beta}A_{r}B}{J_{tot}V_{r}}Q_{rv}(p_{r},u)$
@@
with $K(u)$ defined according to (25) for the case of the real deteriorated valve, and $u$ according to (24).
VII Results
The controllers are tested with the same tuning both in simulation and in the experimental setup. Furthermore, the same gains are chosen for the baseline solution and the presented algorithm, with the latter having an additional tuning parameter $\gamma$ for regulating the adaptation rate in the dynamic inversion. Different solutions for different values for $\gamma$ are presented to analyze its impact.
@@ VII-A Tracking performance
The experimental recordings for the two controllers are presented in Fig. 10, and in Fig. 11 in a zoomed version.
Both the baseline FL and the proposed algorithm are able to track the reference with an error which is always smaller than 0.01 rad. Around 180 seconds tracking is temporary lost or both solutions, due to limitations in the common supply pressure shared by both hydraulic systems. Both controllers saturate and are then able to track the reference back as soon as the required force (hence flow) decreases for both. A quantitative comparison between the two solutions, discarding the interval of saturation, is presented later.
Figure 10: Experimental tracking performance of the Feedback linearization based controller, with and without applying the dynamical mapping inversion.
@@
However, from the zoomed plot it is already possible to appreciate some differences between the two solutions. The control input $u$ displays reduced chattering, especially in the region where no pitching action is present. The control action appears smoother also in the region of active pitching, even if it also contains more similar oscillations with respect to the baseline solution. Such difference is smoothened when analyzing the actual spool’s position $x_{v}$. In fact, the valve performs a natural filtering action according to its bandwidth.
The signal quality in $u$ and $x_{v}$ are equivalent for the MI solution, while part of the noise is filtered out by the valve dynamics in the standard solution, since the controller doesn’t have knowledge of the valve dynamics, nor of valve limitations.
Remark 4.
The spikes in the valve position can be originated by the internal closed-loop or by the internal position sensor.
@@ VII-B Mapping inversion algorithm
The tracking performance of the mapping inversion algorithm is presented in Fig. 12, both for the case of simulation and the experiment. The slight mismatch between the two can be related to model uncertainty, and to the fact that the digitally reproduced inertia is an equivalent load which wouldn’t be present in the real turbine. Its effect on the total load $M_{h,l}$ is the addition of a high-frequency component which causes low-amplitude high-frequency oscillations in the signals, reflected in a virtual input displaying larger peaks. $v$ and $\rho$ are the signal defined in (46), while $v_{sim}$ is the equivalent of the $v$ signal which is obtained in simulation, hence based on the real states of the system and not on the noisy measurements. The algorithm is not able to perfectly reconstruct $v$, because of the presence of noise in the signals.
Figure 12: Dynamical mapping inversion algorithm: Simulation with correct virtual input reconstruction, and experiment.
However, it is shown how it is possible to determine a proper tuning such that it is possible to reconstruct the actual virtual input in the plant ($v_{sim}$) and filtering out the noise present in the measurements. Thus, there is a natural filtering action which can be approximated with a simple first order low-pass filter in order to determine an indication of the bandwidth of such filtering action, in comparison to the bandwidth of the valve which represents a physical limitation in what the control action can deliver on the real system. The comparison, for $\gamma$ values of $(5,\,10,\,20)\times 10^{-4}$, is presented in Fig. 13.
The cutoff frequency of the MI algorithm is always lower than the valve’s bandwidth, meaning the valve only needs to follow trajectories it can actually withstand. With respect to the standard solution, where the filtering was occurring in the valve directly, in this new solution is the algorithm accounting for the valve limitations and filtering the input signal. This result support the similar signal quality in $u$ and $x_{v}$ in the MI solution that was found in Fig. 11.
@@
The tuning presented shows some room with respect to the maximum bandwidth that can be achieved by the valve. However, a faster tuning is reflected in a noisier reconstruction of $v_{sim}$ from the noisy measurements, leading to instability. The slew rate limitation of the valve may also play a role.
Figure 13: Bode plots of valve dynamics and mapping inversion filtering effect for different adaptation gains.
@@ VII-C Controllers comparison
Assessing the tracking performance based on the error is straightforward, while assessing the signal quality of the input is more challenging since it is not a constant average signal. The solution was to calculate a moving average $\bar{u}$ and defining a new signal $u_{f}=u-\bar{u}$ which has zero mean and still maintains the information of the high frequency oscillation. The moving average window is the same for all the tests. The same applies for determining $x_{v,f}$. The case of $u_{f}$, for the same moving average filter applied to both solutions, is illustrated in Fig. 14.
@@
  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{RMS}(u_{f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}u_{fi}^{2}}$
  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{RMS}(x_{v,f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{v,fi}^{2}}$
  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{MAX}(e_{\beta})=\max_{0\leq t\leq T}|e_{\beta}(t)|$
  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{RMS}(e_{\beta})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}e_{\beta,i}^{2}}$
  • $\text{QWAE}(e_{\beta})=\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$
  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{ISE}(e_{\beta})=\int_{0}^{T}e_{\beta}^{2}(t)\,dt$
where $T$ is the simulation time and the resulting values are normalized for having a clearer comparison. The QWAE (Quadratically Weighted Average Error) metric is used to penalize larger errors more heavily, while the ISE (Integral Squared Error) metric measures the persistence of errors. Finally, the time window where all the controllers saturate is removed from the evaluation, since it is not representative of the real turbine operation.
Figure 14: Input signal $u$ and moving average $\bar{u}$ (upper), and the difference $u-\bar{u}$ (lower). Left column plots are with FL alone, right column is with the MI combined with FL. Chattering is significantly reduced with inclusion of MI.
The experimental results for both controllers are presented in Fig. 15, displaying how the solution proposed in the article outperforms the baseline for the same controller tuning and with the adaptation gain $\gamma=10\times 10^{-4}$, which corresponds to the tuning case $MIFL,10$ in Fig. 16. The adaptive law was implemented using forward Euler integration, with a sampling time of 1ms.
Figure 15: Experimental results comparing the baseline algorithm Fig. 2 with the suggested MI approach Fig. 3 . Blue bars of the baseline control are normalized to 1, and green bars (scaled accordingly) show the quality measures for the MIFL solution, relative to that of the baseline approach (FL). Introduction of the the mapping inversion has very significant benefits. Normalization value presented on top of each group of bars.
@@
[…]
Figure 16: Experimental results comparing different tunings for the input inversion algorithm. Higher values of $\gamma$ are reflected in smaller error indexes, but larger chattering amplitude. Normalization values on top of each bar group is added to make comparison possible with Fig.15.
The results for three sets of tuning parameters of the inversion algorithm, see Fig. 13, are presented in Fig. 16. Except for the maximum error metrics, all the others share the expected trend: error values decrease with higher values of $\gamma$, while input chatter values increase. The normalization value presented on top of each bar group is always smaller than the respective one presented in Fig. 15, meaning that the solution with the MI outperforms the baseline FL for all the tuning cases of the MI.
The experimental results hence demonstrate that the MI approach, with the $\gamma$ tuning described, provides an easily implementable and effective solution that can obtain satisfactory tracking of an electro hydraulic actuator while obtaining significant reduction in wear caused by reduction in EHA micro movements and simultaneously by reduced energy in unwanted valve fluctuations.
@@ VIII Conclusion
This article presented a solution to eliminate the formation of an algebraic loop when designing conventional nonlinear controllers. This was achieved by design of a virtual input and dynamically deriving a control action, which emulated the desired behavior, while control valve trajectory commands were avoided if they would exceed the specified admissible range for valve operations.
The effectiveness of the proposed solution was tested in high-fidelity simulation and experimentally for the industrial application of wind turbines’ blade pitch control, showing a general improvement in tracking performance and an essential reduction of input chatter, especially in the regions where the pitch reference is highly varying.
The dynamic inversion solution enabled room for flexibility and allows for future progress with other nonlinear parameter estimation algorithms. A further step towards industrial implementation will be to derive the implicit discrete-time version of the algorithm.
@@ Acknowledgements
This research was funded by The Danish Energy Technology Development and Demonstration Programm (EUDP) through the project: “Decreased Cost of Energy (CoE) from wind turbines by reducing pitch system faults”, grant number 64022-1058. The authors much appreciate this support. The authors would also like to thank Prof. H. C. Pedersen, as well as postdoctoral researchers M. Bhola and G. Wrat, and the students at Aalborg University, for their valuable assistance in conducting the experiments.